# Impact of unmodeled eccentricity on the tidal deformability measurement and implications for gravitational wave physics inference

### v1

@@ I Introduction
The systematic errors are independent of SNR, while the statistical errors scale as the inverse of SNR. Hence, for loud events (high SNR), the systematic errors can easily become greater than the statistical errors. Since CE and ET will observe GW events with higher SNR, leading to a precise measurement of binary parameters like the tidal deformability, and hence the equation of state of NS [69, 70, 71] and the cosmological parameters [72, 73], therefore, systematic errors can easily bias these parameters.
@@ II Waveform model
We consider equal mass binaries with $\hat{\lambda}_{1}=\hat{\lambda}_{2}=\hat{\lambda}$ i.e., $\tilde{\Lambda}\to\hat{\lambda}$ and $\delta\tilde{\Lambda}\to 0$. This leads to the reduction in the number of parameters to be estimated [60]. The contribution of $\delta\tilde{\Lambda}$ (compared to the $\tilde{\Lambda}$) is small and can be neglected. Although the tidal contribution to phase has been calculated to even higher PN orders, Eq. (10) is sufficient for the present context as the aim is to compute the systematic error due to eccentricity on tidal deformability given that the eccentricity is a low-frequency effect.
@@ III Error analysis
where *** denotes complex conjugation, ‘,’ denotes the partial derivative with respect to parameter $\theta_{a}$, $\tilde{h}(f)$ is the Fourier transform of $h(t)$ and $S_{n}(f)$ is the noise power spectral density (PSD) of the detector. The SNR $\rho$ for a signal $h(t)$ is defined as
@@ IV.1 Implications of biased $\tilde{\Lambda}$ for NS EoS
For NS with mass $m_{s}=1.4M_{\odot}$, Eq. (IV.1) predicts $\tilde{\Lambda}=528.28$ which represents the true value of $\tilde{\Lambda}$. Systematic bias on $\tilde{\Lambda}$ and $m_{s}$ introduces a systematic shift in their inferred value, i.e., $\tilde{\Lambda}+\Delta\tilde{\Lambda}$ and $m_{s}+\Delta m_{s}$. The shift in the value of $\tilde{\Lambda}$ and $m_{s}$ can either be positive or negative depending on the sign of the systematic bias, which in turn depends on the correlation of $\tilde{\Lambda}$ and $m_{s}$ with $e_{0}$. The inset plots in Figure 2 show that the systematic errors on $\tilde{\Lambda}$ are negative. In other words, neglecting eccentricity underestimates the value of $\tilde{\Lambda}$.
@@ IV.2 Implications of biased $\tilde{\Lambda}$ for redshift inference via Love siren method
This section analyzes the consequences of the biased $\tilde{\Lambda}$ estimate on the redshift inference via the Love siren method [112, 113, 72, 73]. Unlike the previous analysis, where errors on redshift were neglected, we now calculate the errors on redshift measurement using the Love siren technique. The observations of GW signals provide the detector frame mass $m_{\rm det}$, which is related to the source frame mass $m_{s}$ by the cosmological redshift

### v2

@@ I Introduction
The systematic errors are independent of SNR, while the statistical errors scale as the inverse of SNR. Hence, for loud events (high SNR), the systematic errors can easily become greater than the statistical errors. Since CE and ET will observe GW events with higher SNR, leading to a precise measurement of binary parameters like the tidal deformability, and hence the equation of state of NS [69, 70, 71] and the cosmological parameters [72, 73, 74], therefore, systematic errors can easily bias these parameters.
@@ II Waveform model
We consider equal mass binaries with $\hat{\lambda}_{1}=\hat{\lambda}_{2}=\hat{\lambda}$ i.e., $\tilde{\Lambda}\to\hat{\lambda}$ and $\delta\tilde{\Lambda}\to 0$. This leads to the reduction in the number of parameters to be estimated [60]. The contribution of $\delta\tilde{\Lambda}$ (compared to the $\tilde{\Lambda}$) is small and can be neglected. Although the tidal contribution to phase has been calculated to even higher PN orders [150], Eq. (10) is sufficient for the present context as the aim is to compute the systematic error due to eccentricity on tidal deformability given that the eccentricity is a low-frequency effect.
@@ III Error analysis
where * denotes complex conjugation, ‘,’ denotes the partial derivative with respect to parameter $\theta_{a}$, $\tilde{h}(f)$ is the Fourier transform of $h(t)$ and $S_{n}(f)$ is the noise power spectral density (PSD) of the detector. The SNR $\rho$ for a signal $h(t)$ is defined as
@@ IV.1 Implications of biased $\tilde{\Lambda}$ for NS EoS
For NS with mass $m_{s}=1.4M_{\odot}$, Eq. (IV.1) predicts $\tilde{\Lambda}=528.28$ which represents the true value of $\tilde{\Lambda}$. Systematic bias on $\tilde{\Lambda}$ and $m_{s}$ introduces a systematic shift in their inferred value, i.e., $\tilde{\Lambda}+\Delta\tilde{\Lambda}$ and $m_{s}+\Delta m_{s}$. The shift in the value of $\tilde{\Lambda}$ and $m_{s}$ can either be positive or negative depending on the sign of the systematic bias, which in turn depends on the correlation of $\tilde{\Lambda}$ and $m_{s}$ with $e_{0}$. The inset plots in Figure 2 show that the systematic errors on $\tilde{\Lambda}$ are negative. In other words, neglecting eccentricity underestimates the value of $\tilde{\Lambda}$. This can be qualitatively understood by looking at the number of GW cycles due to eccentricity and tidal deformability. Increasing the magnitude of the tidal deformability parameter leads to an increase in the number of inspiral cycles, while the number of cycles decreases with an increase in eccentricity [60]. An eccentric GW signal has less number of cycles compared to a circular signal, which is also the case for lower tidal deformability. Thus, when a quasicircular template is used for analyzing an eccentric signal, it underestimates the value of tidal deformability.
@@ IV.2 Implications of biased $\tilde{\Lambda}$ for redshift inference via Love siren method
This section analyzes the consequences of the biased $\tilde{\Lambda}$ estimate on the redshift inference via the Love siren method [113, 114, 72, 73, 74]. Unlike the previous analysis, where errors on redshift were neglected, we now calculate the errors on redshift measurement using the Love siren technique. The observations of GW signals provide the detector frame mass $m_{\rm det}$, which is related to the source frame mass $m_{s}$ by the cosmological redshift